# Deep learning framework for action prediction reveals multi-timescale locomotor control

### v4

@@ Methods
To better understand swing foot prediction timescale, we aim to identify two critical timing events: (1) the time at which the swing foot begins its swing phase and (2) the time at which it starts containing significant information about future foot placement. The onset of swing phase is defined as the peak velocity of the swing foot along the fore-aft direction. From this peak, we trace backward to identify the point where the velocity drops below 5% of its maximal value (Figure 5B). This approach of anchoring the swing detection to the peak velocity ensures robustness and allows to avoid false positives.
@@
We analyze the autoregressive baseline $R^{2}$ curve to determine the time at which the swing foot kinematics begins predicting substantially the future foot placement. The $R^{2}$ curve typically shows a non-decreasing trend as the gait cycle progresses, eventually reaching $R^{2}=1$ at the final timestep. A notable “breakpoint” in the $R^{2}$ curve is the phase at which the predictive power of the swing foot kinematics sharply rises. To systematically identify this breakpoint, we calculate the average slope of the $R^{2}$ curve, defined as $r=R^{2}(1)-R^{2}(0)$, and locate the gait phase where the slope reaches its maximal value. We then trace backward from this gait phase to find the first point where the slope falls below the average slope, marking the critical transition phase (Figure 5C). This approach provides a robust method for identifying key transitions in the gait cycle and aligns with the underlying dynamics of gait and motor control.
@@ Acknowledgments
  • FCNN: the fully connected layers are designed such that the number of nodes in each layer decays exponentially, with a minimum of 8 nodes per layer and a guaranteed minimum number of layers. The decay factor, decay$\in\{2,4,8,16\}$; the dropout rate, dropout$\in\{0,0.1,0.2,0.3\}$.
@@
Figure 6: Predictive power across models. A Predictive power during treadmill walking. The vertical line indicates the critical phase when the relative predictive power is maximized, thereby only gait phases before the gray-shaded area are considered. The model with the average highest score prior to the critical phase was deemed optimal, and its predictive power is shown by the interpolated curve (LOWESS). See the Methods section for more details on the evaluation metric. The normalized model score is summarized in Table 1. B Predictive power during treadmill running.

### v5

@@ Methods
To better understand swing foot prediction timescale, we aim to identify two critical timing events: (1) the time at which the swing foot begins its swing phase and (2) the time at which it starts containing significant information about future foot placement. The onset of the swing phase is defined as the peak velocity of the swing foot along the fore-aft direction. From this peak, we trace backward to identify the point where the velocity drops below 5% of its maximal value (Figure 6B). This approach of anchoring the swing detection to the peak velocity ensures robustness and allows to avoid false positives.
@@
We analyze the autoregressive baseline $R^{2}$ curve to determine the time at which the swing foot kinematics begins to substantially predict future foot placement. The $R^{2}$ curve typically shows a non-decreasing trend as the gait cycle progresses, eventually reaching $R^{2}=1$ at the final timestep. A notable “breakpoint” in the $R^{2}$ curve is the phase at which the predictive power of the swing foot kinematics sharply rises. To systematically identify this breakpoint, we calculate the average slope of the $R^{2}$ curve, defined as $r=R^{2}(1)-R^{2}(0)$, and locate the gait phase where the slope reaches its maximal value. We then trace backward from this gait phase to find the first point where the slope falls below the average slope, marking the critical transition phase (Figure 6C). This approach provides a robust method for identifying key transitions in the gait cycle and aligns with the underlying dynamics of gait and motor control.
@@ Acknowledgments
[…]
@@
Figure 7: Predictive power across models. A Predictive power during treadmill walking. The vertical line indicates the critical phase when the relative predictive power is maximized, thereby only gait phases before the gray-shaded area are considered. The model with the average highest score prior to the critical phase was deemed optimal, and its predictive power is shown by the interpolated curve (LOWESS[78]). See the Methods section for more details on the evaluation metric. The normalized model score is summarized in Table 1. B Predictive power during treadmill running.